Write an equation of the line with the given properties. Your answer should be written in standard form.
m = −3 passing through P(2, 8)

Answer by Cromlix(4381)

using the form y - b = m(x - a)
y - 8 = -3(x - 2)
y - 8 = -3x + 6
y = -3x + 6 + 8
y = -3x + 14
Hope this helps.
:-)